OpenStudy (christos):

Please help me solve 0 < (2x-4)^-1< 1/2

OpenStudy (anonymous):

\[0< \frac{ 1 }{ 2 x-4 }<\frac{ 1 }{ 2 }\] \[x>3\]

OpenStudy (anonymous):

oh? wow that is a great website? yeah does it include step by step though?

hartnn (hartnn):

yes. but it just gives solution, does not tutor...

OpenStudy (anonymous):

oh :(

OpenStudy (christos):

How did you get from step 1 to step 2 :S

OpenStudy (mrhoola):

are you still having problems ? or is this solved ?

OpenStudy (christos):

I mean I cant understand how he got from step 1 to step 2 ( first poster)

hartnn (hartnn):

if a/b < c/d then b/a > d/c so, when 1/ (2x-4) < 1/2 then , 2x-4 > 2

OpenStudy (christos):

Got it now! thanks!

OpenStudy (mrhoola):

goku saved the day !

OpenStudy (christos):

:P

hartnn (hartnn):

:)

OpenStudy (christos):

One last question am I able to remove the 0 < part because it is zero?

hartnn (hartnn):

no, since we got x>3 means 'x' is positive, now if we multiply 2x-4 on all side, we get , 0< 1 which is obviously true, but will not give us any inequality in x so, only inequality we have is x>3

OpenStudy (mrhoola):

x>3 <--- kinda looks like a chicken/turkey squinting his/her eyes

hartnn (hartnn):

then what does the inequality i <3u look like ? :P

OpenStudy (christos):

So this is what I get to if I try to solve it by my own 0 < 1/(2x-4)^-1<1/2 0<1/2x-4<1/2 0<2x-4<2 -4<2x<-2 -2<x<-1 Where is my mistake?

OpenStudy (mrhoola):

im flattered LOL jk

OpenStudy (mrhoola):

third line cristo !!

hartnn (hartnn):

flip the sign!

OpenStudy (christos):

2nd line was mistyped it should be 0 < 1/2x-4<1/2

OpenStudy (christos):

hmm sec

OpenStudy (mrhoola):

(2x-4)*0 =0 (2x-4)*1/2x-4 =1 and (2x-4)*1/2 =x-2

OpenStudy (christos):

On which line do I need to flip the sign ? :D

OpenStudy (christos):

and why? :D

hartnn (hartnn):

because of this : if a/b < c/d then b/a > d/c so, 1/(2x-4) < 1/2 gives you 2x-4 > 2

hartnn (hartnn):

don't consider both inequalities together, treat 0 < 1/(2x-4) and 1/(2x-4) < 1/2 separately.

OpenStudy (christos):

I solved it! Thank you all!!!!!!!!

hartnn (hartnn):

welcome :)